# A unified convention for achievement positional games

### v1

@@ 4.1 AchievementPos(2,2) is in P
We start by showing that the game is easily solved when all edges have size at most 2.
@@
AchievementPos(2,2) is in P. More precisely, there exists an algorithm which, given an achievement positional game $\mathcal{G}=(V,E_{L},E_{R})$ such that all elements of $E_{L}\cup E_{R}$ have size at most 2, decides whether Left has a winning strategy on $\mathcal{G}$ as first (resp. second) player in time $O(|V|^{2}|E_{L}|+|V||E_{R}|)$.
@@
First of all, let us explain how to reduce to the case where all elements of $E_{L}\cup E_{R}$ are of size exactly 2. Let $\alpha$ (resp. $\beta$) be the number of edges of size 1 whose color corresponds to the first (resp. second) player. The case $\alpha\geq 1$ is trivially winning for the first player, while the case where $\alpha=0$ and $\beta\geq 2$ is trivially winning for the second player. In the remaining case $(\alpha,\beta)=(0,1)$, the first player is forced to pick a particular vertex $u$, so we may replace $\mathcal{G}$ with $\mathcal{G}_{u}$ (resp. $\mathcal{G}^{u}$) if Left (resp. Right) starts, which can be computed in time $O(|E_{L}|+|E_{R}|)$. We can repeat this operation as many times as needed, say $t$ times, until all edges have size exactly 2. Up to this preprocessing step in time $O(|V|(|E_{L}|+|E_{R}|))$, we thus assume that we have a blue graph $G_{L}=(V,E_{L})$ and a red graph $G_{R}=(V,E_{R})$.
Furthermore, we can assume that Right is now the first player, i.e. $t$ is odd if Left started the game or even if Right started the game. Indeed, the case where Left is next to play is straightforward: if there is a blue $P_{3}$ (path on 3 vertices), then Left can win in two moves, otherwise Right has a non-losing pairing strategy by Lemma 3.5. Finally, we can assume that there is no red $P_{3}$ i.e. the red edges form a matching, otherwise Right can win in two moves. Therefore, Right has a non-losing strategy if and only if she can avoid any situation where it is Left’s turn and there is an intact blue $P_{3}$ $uvw$ such that Left can pick $v$ without losing on the spot. As such, Right wants to force all of Left’s moves and keep the initiative until she has picked at least one vertex in each blue $P_{3}$. If she manages that, then the updated blue edges will form a matching, giving her a non-losing pairing strategy by Lemma 3.5. We now go on with a rigorous proof.
Given a vertex $u$, an alternating $u$-path is a path $u_{1}\cdots u_{s}$ in $(V,E_{L}\cup E_{R})$ such that: $u_{1}=u$, $u_{i}u_{i+1}\in E_{R}$ for all odd $11\leq i\leq s-$ 1, and $u_{i}u_{i+1}\in E_{L}$ for all even $11\leq i\leq s-$ 1. We accept the case $s=1$ in this definition. We split all possibilities for Right’s next move $u$ into three types, as follows. We say $u$ is of type (1) if there exist an alternating $u$-path $P=u_{1}\cdots u_{s}$ with odd $s\geq 3$ and a vertex $y\not\in P$ such that $u_{s-1}v\in E_{L}$. Otherwise, there exists a unique maximal alternating $u$-path, which we denote by $P(u)=u_{1}\cdots u_{s}$. We then say $u$ is of type (2) if $s$ is odd, or of type (3) if $s$ is even. Note that each $e\in E_{L}$ containing some $u_{i}$ with even $i$ also contains some $u_{j}$ with odd $j$. See Figure 6.
@@
  • We claim that picking $u$ of type (2) is a losing move for Right if and only if there exists a $P_{3}$ in $G_{L}-P(u)$, which denotes the graph obtained from $G_{L}$ by deleting all vertices of $P(u)=u_{1}\cdots u_{s}$. Indeed, suppose Right picks $u_{1}$, which forces Left to pick $u_{2}$, and so on until Right eventually picks $u_{s}$ (recall that $s$ is odd). At this point, since each $u_{i}$ with odd $i$ has been picked by Right, there is no blue edge in which Left has picked a vertex and Right has not picked the other. Moreover, we know $u_{s}$ is isolated in $G_{R}$ by maximality of $P(u)$, so Left has no forced move. If there exists a $P_{3}$ $uvw$ in $G_{L}-P(u)$, then Left can pick $v$ and win the game with his next move. Otherwise, Right can apply a non-losing pairing strategy.
  • We claim that picking any $u$ of type (3) is an optimal move for Right, be it losing or non-losing. Indeed, write $P(u)=u_{1}\cdots u_{s}$ (recall that $s$ is even). We want to show that the situation after the sequence of forced moves triggered by $u=u_{1}$ is not worse for Right than it was before. In other words, we want to show that if Left has a winning strategy $\mathcal{S}$ on $\mathcal{G}^{u_{1},u_{3},\ldots,u_{s-1}}_{u_{2},u_{4},\ldots,u_{s}}=(V^{\prime},E^{\prime}_{L},E^{\prime}_{R})$ as second player then he also has a winning strategy on $\mathcal{G}$ as second player. Such a strategy may be derived from $\mathcal{S}$ as follows. Whenever Right picks some $u_{i}$, Left picks its unique red neighbor (forced move). Otherwise, Left answers according to $\mathcal{S}$. By assumption on $\mathcal{S}$, Left will fill some $e\in E_{L}^{\prime}\subseteq E_{L}$ before Right can fill some $e^{\prime}\in E_{R}^{\prime}$, thus winning the game $\mathcal{G}$ at the same time.
@@
The outcome is characterized by the answers to each of the four combinations of the following question: “Does Left/Right have a winning strategy as first/second player?” By Theorem 4.1, each is solved in time $O(|V|^{2}(|E_{L}|+|E_{R}|))$. ∎
@@ 4.2 AchievementPos(2,3) is coNP-complete
AchievementPos($2$,$3$) is coNP-complete.
@@
We consider the complement of this problem, or rather an equivalent version of it. We show that it is NP-complete to decide whether Left has a non-losing strategy as first player on an achievement positional game where all blue (resp. red) edges have size at most 3 (resp. at most 2).
Let us first show membership in NP. Consider the following strategy $\mathcal{S}$ for Right: pick some $u$ that wins in one move if such $u$ exists, otherwise pick some $v$ that prevents Left from winning in one move if such $v$ exists, otherwise pick some $w$ at the center of an intact red $P_{3}$ is one exists, otherwise (Left has a non-losing pairing strategy) pick an arbitrary vertex. Clearly, if Right has a winning strategy on $\mathcal{G}$ as second player, then $\mathcal{S}$ is one. Moreover, the move prescribed by $\mathcal{S}$ in any given situation is easily computed in polynomial time. Therefore, a polynomial certificate for Left’s non-losing strategy is simply the sequence of all of Left’s moves, assuming that Right plays according to $\mathcal{S}$.
We now reduce 3-SAT to our problem. Consider an instance $\phi$ of 3-SAT, with a set of variables $V$ and a set of clauses $C$. We build a game $\mathcal{G}=(V,E_{L},E_{R})$ as follows (see Figure 7 for a visual example):
@@
  • $V=\bigcup_{x\in V}V_{x}\cup\bigcup_{c\in C}V_{c}\cup\left\{\omega,\check{\omega},\hat{\omega}\right\}$.
  • $E_{L}=\bigcup_{x\in V}\left\{\left\{x,\neg x\right\}\right\}\cup\bigcup_{c=\ell_{1}\vee\ell_{2}\vee\ell_{3}\in C}\left\{\left\{\ell_{1},c_{\ell_{1}},c^{\prime}_{\ell_{1}}\right\},\left\{\ell_{2},c_{\ell_{2}},c^{\prime}_{\ell_{2}}\right\},\left\{\ell_{3},c_{\ell_{3}},c^{\prime}_{\ell_{3}}\right\}\right\}$.
  • $E_{R}=\bigcup_{c=\ell_{1}\vee\ell_{2}\vee\ell_{3}\in C}\left\{\left\{c_{\ell_{1}},c_{\ell_{2}}\right\},\left\{c_{\ell_{2}},c_{\ell_{3}}\right\},\left\{c_{\ell_{3}},c_{\ell_{1}}\right\}\right\}\cup\left\{\left\{\omega,\check{\omega}\right\},\left\{\omega,\hat{\omega}\right\}\right\}$.
@@
Since there are multiple pairwise vertex-disjoint red $P_{3}$’s, every move from Left must threaten to win on the next move until the last red $P_{3}$ is broken.
As such, Left must start by picking $\ell\in\{x,\neg x\}$ for some $x\in V$, which forces Right to pick the other one since $\{x,\neg x\}\in E_{R}$. In the updated game, for each clause $c$ containing the literal $\ell$, we get the blue edge $\{c_{\ell},c^{\prime}_{\ell}\}$, which is the only edge containing $c^{\prime}_{\ell}$. Therefore, Lemma 3.8 ensures that it is optimal for Left to pick $c_{\ell}$ for each clause $c$ which contains $\ell$, which forces Right to pick $c^{\prime}_{\ell}$ in response. This breaks every red $P_{3}$ in the clause gadgets corresponding to clauses containing $\ell$.
Left must repeat this process of picking a literal and then breaking all clause gadgets of clauses containing that literal, until he has picked at least one of $c_{\ell_{1}}$, $c_{\ell_{2}}$ or $c_{\ell_{3}}$ for each clause $c=\ell_{1}\vee\ell_{2}\vee\ell_{3}\in C$.
If there exists a valuation $\mu$ which satisfies $\phi$, then Left succeeds in doing so, by picking $x$ if $\mu(x)={\sf T}$ or $\neg x$ if $\mu(x)={\sf F}$, for all $x\in V$. After that, he can simply pick $\omega$, thus ensuring not to lose the game. If such a valuation does not exist, then Left will have to play a move that does not force Right’s answer while leaving at least one red $P_{3}$ intact, thus losing the game. All in all, Left has a non-losing strategy on $\mathcal{G}$ as first player if and only if $\phi$ is satisfiable, which concludes the proof. ∎
@@
In the previous construction, Left never has a winning strategy as first player. Therefore, it is NP-hard to decide whether optimal play leads to a draw when Left starts on an achievement positional game with blue edges of size at most 3 and red edges of size at most 2.
4.3 AchievementPos($3$,$2$) is NP-hard
@@
We use the same construction as in the previous proof (see Figure 7), except we add two copies of the blue butterfly gadget (recall Figure 1) to transform draws into wins for Left.
Left cannot play on any of the two butterflies as long as he has not broken every red $P_{3}$, as Right would then win in two moves. As such, he must play the same way as in the previous proof until every red $P_{3}$ is broken. Since Right’s moves are all forced during that phase of play, she cannot play on a butterfly either. If $\phi$ is not satisfiable, then Right wins. Otherwise, Left breaks the last red $P_{3}$ by picking $\omega$, and then Right can only break one of the two blue butterflies. Left plays first on the other butterfly, thus winning the game in two more moves. ∎
@@ 4.4 AchievementPos(3,3) is PSPACE-complete
The classic 3-QBF decision problem can be formulated in terms of the following game. The input is a logic formula $\phi$ in CNF form, with clauses of size exactly 3 and an even number of variables which have a prescribed ordering $x_{1},\ldots,x_{2n}$. Two players, Satisfier and Falsifier, take turns choosing valuations for the variables. Satisfier goes first, setting $x_{1}$ to T or F. Falsifier then sets $x_{2}$ to T or F, and so on until a full valuation $\mu$ is built. Note that Satisfier (resp. Falsifier) takes care of all variables whose index is odd (resp. even). Satisfier wins the game if $\mu$ satisfies $\phi$, otherwise Falsifier wins.

### v2

@@ 4.1 AchievementPos(2,2) is in P
We start by showing that the game is easily solved when all edges have size at most 2. The following definition will be useful here and also later in the paper. A blue $P_{3}$ (resp. red $P_{3}$) is a pair of blue (resp. red) edges of the form $\{\{u,v\},\{v,w\}\}$, which may be denoted as uvwuvw. Note that a blue (resp. red) $P_{3}$ threatens a two-move win for Left (resp. Right) by picking $v$ and then picking $u$ or $w$.
@@
AchievementPos(2,2) is in P. More precisely, there exists an algorithm which, given an achievement positional game $\mathcal{G}=(V,E_{L},E_{R})$ such that all elements of $E_{L}\cup E_{R}$ have size at most 2, decides whether Left has a winning strategy on $\mathcal{G}$ as the first (resp. second) player in time $O(|V|^{2}|E_{L}|+|V||E_{R}|)$.
@@
First of all, let us explain how to reduce to the case where all elements of $E_{L}\cup E_{R}$ have size exactly 2. Let $\alpha$ (resp. $\beta$) be the number of 1-edges whose color corresponds to the first (resp. second) player. The case $\alpha\geq 1$ is trivially winning for the first player, while the case where $\alpha=0$ and $\beta\geq 2$ is trivially winning for the second player. In the remaining case $(\alpha,\beta)=(0,1)$, the first player is forced to pick a particular vertex $u$, so we may replace $\mathcal{G}$ with $\mathcal{G}_{u}$ (resp. $\mathcal{G}^{u}$) if Left (resp. Right) starts, which can be computed in time $O(|E_{L}|+|E_{R}|)$. We can repeat this operation as many times as needed, say $t$ times, until all edges have size exactly 2. Up to this preprocessing step in time $O(|V|(|E_{L}|+|E_{R}|))$, we thus assume that we have a blue graph $G_{L}=(V,E_{L})$ and a red graph $G_{R}=(V,E_{R})$.
Furthermore, we can assume that Right is now the first player, i.e. $t$ is odd if Left started the game or even if Right started the game. Indeed, the case where Left is next to play is straightforward: if there is a blue, then Left can win in two moves, otherwise Right has a non-losing pairing strategy by Lemma 3.5. Finally, we can assume that there is no red $P_{3}$ i.e. the red edges form a matching, otherwise Right can win in two moves. Therefore, Right has a non-losing strategy if and only if she can avoid any situation where it is Left’s turn and there is an intact blue $P_{3}$ uvwuvw such that Left can pick $v$ without losing on the spot. As such, Right wants to force all of Left’s moves and keep the initiative until she has picked at least one vertex in each blue $P_{3}$. If she manages that, then the updated blue edges will form a matching, giving her a non-losing pairing strategy by Lemma 3.5. We now go on with a rigorous proof.
Given a vertex $u$, an alternating $u$-path is a path $u_{1}\cdots u_{s}$ in $(V,E_{L}\cup E_{R})$ such that: $u_{1}=u$, $u_{i}u_{i+1}\in E_{R}$ for all odd $1\leq i\leq s-1$, and $u_{i}u_{i+1}\in E_{L}$ for all even $1\leq i\leq s-1$. We accept the case $s=1$ in this definition. We split all possibilities for Right’s next move $u$ into three types, as follows. We say $u$ is of type (1) if there exist an alternating $u$-path $P=u_{1}\cdots u_{s}$ with odd $s\geq 3$ and a vertex $y\not\in P$ such that $u_{s-1}v\in E_{L}$. Otherwise, there exists a unique maximal alternating $u$-path, which we denote by $P(u)=u_{1}\cdots u_{s}$. We then say $u$ is of type (2) if $s$ is odd, or of type (3) if $s$ is even. Note that each $e\in E_{L}$ containing some $u_{i}$ with even $i$ also contains some $u_{j}$ with odd $j$. See Figure 6.
@@
  • We claim that picking $u$ of type (2) is a losing move for Right if and only if there exists a $P_{3}$ in $G_{L}-P(u)$, which denotes the graph obtained from $G_{L}$ by deleting all vertices of $P(u)=u_{1}\cdots u_{s}$. Indeed, suppose Right picks $u_{1}$, which forces Left to pick $u_{2}$, and so on until Right eventually picks $u_{s}$ (recall that $s$ is odd). At this point, since each $u_{i}$ with odd $i$ has been picked by Right, there is no blue edge in which Left has picked a vertex and Right has not picked the other. Moreover, we know $u_{s}$ is isolated in $G_{R}$ by maximality of $P(u)$, so Left has no forced move. If there exists a $P_{3}$ uvwuvw in $G_{L}-P(u)$, then Left can pick $v$ and win the game with his next move. Otherwise, Right can apply a non-losing pairing strategy.
  • We claim that picking any $u$ of type (3) is an optimal move for Right, be it losing or non-losing. Indeed, write $P(u)=u_{1}\cdots u_{s}$ (recall that $s$ is even). We want to show that the situation after the sequence of forced moves triggered by $u=u_{1}$ is not worse for Right than it was before. In other words, we want to show that if Left has a winning strategy $\mathcal{S}$ on $\mathcal{G}^{u_{1},u_{3},\ldots,u_{s-1}}_{u_{2},u_{4},\ldots,u_{s}}=(V^{\prime},E^{\prime}_{L},E^{\prime}_{R})$ as the second player then he also has a winning strategy on $\mathcal{G}$ as the second player. Such a strategy may be derived from $\mathcal{S}$ as follows. Whenever Right picks some $u_{i}$, Left picks its unique red neighbor (forced move). Otherwise, Left answers according to $\mathcal{S}$. By assumption on $\mathcal{S}$, Left will fill some $e\in E_{L}^{\prime}\subseteq E_{L}$ before Right can fill some $e^{\prime}\in E_{R}^{\prime}$, thus winning the game $\mathcal{G}$ at the same time.
@@
The outcome is characterized by the answers to each of the four combinations of the following question: “Does Left/Right have a winning strategy as the first/second player?” By Theorem 4.1, each is solved in time $O(|V|^{2}(|E_{L}|+|E_{R}|))$. ∎
@@ 4.2 AchievementPos(2,3) is coNP-complete
Proposition 4.3.
AchievementPos(22,$q$) is in coNP for all $q\geq 0$.
Proof of Proposition 4.3.
We show that the complement problem, which asks whether Right has a non-losing strategy as the second player, is in NP for all $q\geq 0$. Consider the strategy $\mathcal{S}$ for Left where, in each turn, Left picks his vertex according to the following order of priority:
  1. 1.
    Pick some $u$ that wins in one move if such $u$ exists.
  2. 2.
    Pick some $v$ that prevents Right from winning in one move if such $v$ exists.
  3. 3.
    Pick some $w$ at the center of an intact blue $P_{3}$ is one exists.
  4. 4.
    (Left has a non-losing pairing strategy) Pick an arbitrary vertex.
Since all blue edges have size at most 2, if Left has a winning strategy on $\mathcal{G}$ as the first player, then $\mathcal{S}$ clearly is one. Moreover, given any possible move for Left, it is easy to check in polynomial time whether that move abides by $\mathcal{S}$ or not. Therefore, a polynomial certificate for Right’s non-losing strategy is simply a sequence of all of moves from both players in which Left plays according to $\mathcal{S}$ and does not win. ∎
@@
AchievementPos(22,$q$) is coNP-complete for all $q\geq 3$.
@@
Membership in coNP is given by Proposition 4.3. As for coNP-hardness, addressing the case $q=3$ is sufficient. We reduce 3-SAT to the complement of AchievementPos(22,33), i.e. the problem of deciding whether Right has a non-losing strategy as the second player on an achievement positional game $\mathcal{G}$ where all blue (resp. red) edges have size at most $2$ (resp. at most $3$). Consider an instance $\phi$ of 3-SAT, with a set of variables $V$ and a set of clauses $C$. We build a game $\mathcal{G}=(V,E_{L},E_{R})$ as follows (see Figure 7 for a visual example):
@@
  • $V=\bigcup_{x\in V}V_{x}\cup\bigcup_{c\in C}V_{c}\cup\left\{\gamma,\omega,\omega_{1},\omega_{2}\right\}$.
  • $E_{L}=\bigcup_{c=\ell_{1}\vee\ell_{2}\vee\ell_{3}\in C}\left\{\left\{c_{\ell_{1}},c_{\ell_{2}}\right\},\left\{c_{\ell_{2}},c_{\ell_{3}}\right\},\left\{c_{\ell_{3}},c_{\ell_{1}}\right\}\right\}\cup\left\{\left\{\omega,\omega_{1}\right\},\left\{\omega,\omega_{2}\right\}\right\}$.
  • $E_{R}=\{\{\gamma\}\}\cup\bigcup_{x\in V}\left\{\left\{x,\neg x\right\}\right\}\cup\bigcup_{c=\ell_{1}\vee\ell_{2}\vee\ell_{3}\in C}\left\{\left\{\ell_{1},c_{\ell_{1}},c^{\prime}_{\ell_{1}}\right\},\left\{\ell_{2},c_{\ell_{2}},c^{\prime}_{\ell_{2}}\right\},\left\{\ell_{3},c_{\ell_{3}},c^{\prime}_{\ell_{3}}\right\}\right\}$.
@@
As the first player, Left obviously needs to start by picking $\gamma$ because of the red edge $\{\gamma\}$. Now, Right is next to play. Since there are multiple pairwise vertex-disjoint blue $P_{3}$’s, every move from Right must threaten to win on the next move until the last blue $P_{3}$ is destroyed. Indeed, Left would otherwise win in two moves by picking the center of a blue $P_{3}$.
Therefore, Right must pick $\ell\in\{x,\neg x\}$ for some $x\in V$, which forces Left to pick the other one since $\{x,\neg x\}\in E_{R}$. In the updated game, for each clause $c$ containing the literal $\ell$, we get the red edge $\{c_{\ell},c^{\prime}_{\ell}\}$, which is the only edge containing $c^{\prime}_{\ell}$. Therefore, Lemma 3.8 ensures that it is optimal for Right to pick $c_{\ell}$ for each clause $c$ which contains $\ell$, which forces Left to pick $c^{\prime}_{\ell}$ in response. This destroys every blue $P_{3}$ in the clause gadgets corresponding to clauses containing $\ell$.
Right must repeat this process of picking a literal and then destroying all clause gadgets of clauses containing that literal, until she has picked at least one of $c_{\ell_{1}}$, $c_{\ell_{2}}$ or $c_{\ell_{3}}$ for each clause $c=\ell_{1}\vee\ell_{2}\vee\ell_{3}\in C$.
If there exists a valuation $\mu$ of the variables which satisfies $\phi$, then Right succeeds in destroying every blue $P_{3}$ in the clause gadgets, by picking $x$ if $\mu(x)={\sf T}$ or $\neg x$ if $\mu(x)={\sf F}$, for all $x\in V$. After that, she can simply pick $\omega$, destroying the final blue $P_{3}$ and thus ensuring not to lose the game thanks to a pairing strategy on the remaining blue edges. If such a valuation does not exist, then Right will have to play a move that does not force Left’s answer and leaves at least one blue $P_{3}$ intact, thus losing the game. All in all, Right has a non-losing strategy on $\mathcal{G}$ as the second player if and only if $\phi$ is satisfiable, which concludes the proof. ∎
@@
In the previous construction, Right never has a winning strategy as the second player. As such, it is NP-hard to decide whether optimal play leads to a draw when Left is the first player on an achievement positional game with blue edges of size at most 2 and red edges of size at most 3. Therefore, by reversing the colors and removing the vertex $\gamma$, we get a construction that shows NP-hardness of deciding whether optimal play leads to a draw when Left is the first player on an achievement positional game with blue edges of size at most 3 and red edges of size at most 2.
4.3 AchievementPos(33,22) is NP-hard
@@
We use the same construction as in Remark 4.5, except that we add two copies of the blue butterfly gadget (recall Figure 1) to transform draws into wins for Left. See Figure 8.
Left cannot play on any of the two butterflies as long as he has not destroyed every red $P_{3}$, as Right would then win in two moves. As such, he must play the same way as Right did in the proof of Theorem 4.4 (with colors reversed), until every red $P_{3}$ is destroyed. Since Right’s moves are all forced during that phase of play, she cannot play on a butterfly either in the meantime. If $\phi$ is not satisfiable, then Right wins. Otherwise, Left destroys the last red $P_{3}$ by picking $\omega$, and then Right can only destroy one of the two blue butterflies. Left is next to play, and wins in three moves using the intact blue butterfly. ∎
@@ 4.4 AchievementPos(3,3) is PSPACE-complete
To establish PSPACE-completeness for AchievementPos(3,3), we actually prove a stronger result:
Theorem 4.7.
Deciding whether Left has a winning strategy as the first player on an achievement positional game with edges of size $2$ or $3$ and the same edges of size $3$ for both colors is PSPACE-complete.
We first remark that, assuming Theorem 4.7 holds, we obtain PSPACE-completeness for intermediate positions of the Maker-Maker game on 3-uniform hypergraphs as a corollary. Note that, since we consider played moves that were not necessarily optimal, this does not imply any hardness result for MakerMaker($3$).
Corollary 4.8.
Deciding whether the first player has a winning strategy for the Maker-Maker game on a 3-uniform hypergraph with one round (one move of each player) having already been played is PSPACE-complete.
Proof.
Let $\mathcal{G}=(V,E_{2,L}\cup E_{3},E_{2,R}\cup E_{3})$ be an achievement positional game where all edges in $E_{2,L}\cup E_{2,R}$ have size $2$ and all edges in $E_{3}$ have size $3$ (this case is PSPACE-complete by Theorem 4.7). We create two new vertices $u_{L}$ and $u_{R}$, and we define the hypergraph $H=(V\cup\{u_{L},u_{R}\},E^{\prime}_{L}\cup E^{\prime}_{R}\cup E_{3})$ where $E^{\prime}_{L}=\{e\cup\{u_{L}\}\mid e\in E_{2,L}\}$ and $E^{\prime}_{R}=\{e\cup\{u_{R}\}\mid e\in E_{2,R}\}$. Note that $H$ is 3-uniform. Consider the Maker-Maker game played on $H$, after one round of play where the first player FP has picked $u_{L}$ and the second player SP has picked $u_{R}$. Any edge $e\in E^{\prime}_{L}$ now behaves like an edge $e\setminus\{u_{L}\}$ except that SP filling that edge would not count (“blue” edge). Similarly, any edge $e\in E^{\prime}_{R}$ now behaves like an edge $e\setminus\{u_{R}\}$ except that FP filling that edge would not count (“red” edge). As for the edges in $E_{3}$, they are untouched and still there for both players to fill. The situation thus reduces to $\mathcal{G}$ with Left playing first, identifying Left with FP and Right with SP. ∎
The rest of this subsection is dedicated to the proof of Theorem 4.7. Note that membership in PSPACE is given by Proposition 2.1. As for PSPACE-hardness, we perform a reduction from the classic 3-QBF decision problem which has been shown PSPACE-complete by Stockmeyer and Meyer [20], or rather from its complement which is also PSPACE-complete since PSPACE=coPSPACE [14].
The problem 3-QBF can be formulated in terms of the following game. A logic formula $\phi$ in CNF form with clauses of size exactly $3$ is given, along with a prescribed ordering of its variables $x_{1},y_{1},\ldots,x_{n},y_{n}$. Two players, Falsifier and Satisfier, take turns assigning truth values to the variables. We assume that Falsifier goes first, setting $x_{1}$ to T or F. Satisfier then sets $y_{1}$ to T or F. Falsifier then sets $x_{2}$ to T or F, and so on until a full valuation $\mu$ is built. Satisfier wins the game if $\mu$ satisfies $\phi$, otherwise Falsifier wins.